OpenStudy (yamyam70):

Verify the Identity. PLease help :) cos x sin x + 2 cos^2 x - sin x -2 cos x ____________________________________ = cos x -1 sin x + 2 cos x

OpenStudy (anonymous):

in the numerator from first two take common cos x and from last two -1 and make factors.

OpenStudy (yamyam70):

is sine the common factor? is this factoring by grouping?

OpenStudy (anonymous):

cosx (sin x+2 cos x) similarly for second

OpenStudy (anonymous):

\[ \cos x \sin x + 2 \cos^2 x - \sin x -2 \cos x = \cos x(\sin x+2\cos x)-\sin x -2\cos x \]

OpenStudy (anonymous):

\[ =\cos x(\sin x +2\cos x)-(\sin x +2\cos x) = (\cos x -1)(\sin x +2\cos x) \]

OpenStudy (anonymous):

The numerator and denominator cancel out.

OpenStudy (anonymous):

This identity only holds true when: \[ \sin x +2\cos x \neq 0 \]

OpenStudy (yamyam70):

I understand now, Thank you :)